### v1

@@ 3.1 Quantum Average Pooling Implementation
A classical average pooling layer consists of a pooling window of size $D\times D$ pixels that passes over an image consisting of $N\times N$ pixels, with a certain stride $K_{\text{STRIDE}}$ RawatCNN ; SHARMA2018377 ; SaffarCNN . We shall focus on the case where the pooling window moves across the image one pixel at a time, corresponding to a stride value $K_{\text{STRIDE}}=1$. For each position of the pooling window, all pixels within the pooling window are averaged together, and this average is output as the value of a pixel in a new image. In classical techniques depending on the size of $K_{\text{STRIDE}}$ and $D$ the number of output pixels will be some number less than $N^{2}$ and hence the pooling corresponds to reducing the dimension of the image. In the quantum technique, there is no penalty for calculating all $N^{2}$ terms as the operations are done in parallel on the quantum state, hence we shall consider this case and leave the dimensionality reduction as a subroutine that can be implemented after the averaging.
@@
so that they can be added to the average. The pixels accessed by $T_{\Delta x,\Delta y}$ define the pooling tile. For occasions in which the pooling window covers pixels outside the image, it is common to include some padding of the image Yu2023EfficientMP , for example, any pixels outside the $N\times N$ region can be padded by zeros, giving a padded image of dimension $(N+2D)\times(N+2D)$.
There have been many investigations focused on finding quantum parallels to CNN models, whereby measurements are taken of certain qubits such that the dimensionality of the data is reduced. For example, in some setups the pooling layer corresponds to a variational ansatz followed by a measurement of a qubit Hur_2022 . In Cong_2019 an effective measurement was chosen corresponding to the symmetry and details of their specific problem concerning Quantum Phase Estimation. Although these popularised methods achieve the goal of pooling through reducing the dimensionality of the quantum state, they have not focused on giving a replication of an averaging pooling layer from classical CNN models for image data. In this section, we shall therefore demonstrate how LCU methods can be used to perform average pooling on quantum encoded image data. We shall utilise amplitude encoding with distinct coordinates $x$ and $y$ for the image, although any encoding for the data could be used, as long as the correct transformations are implemented with the LCU method.
@@
where $\Omega^{2}=\sum v_{i,j}^{2}$ is a normalisation constant and $v_{i,j}\in\mathbb{R}$. Once we have an image encoded into a quantum state in this manner, we can consider the type of operation required by average pooling. In a simple case, we can consider only the $x$ register for $D=2$. We will need to apply an operation that takes $\lvert x\rangle\rightarrow\frac{1}{\Omega^{\prime}}(\lvert x\rangle+\lvert x\ominus 1\rangle)$ for some normalisation constant $\Omega^{\prime}$, to perform an averaging over pixels amplitudes. This will result in the state $\lvert x\rangle$ having its amplitude transformed to the sum of the amplitudes of the $\lvert x\rangle$ and $\lvert x\oplus 1\rangle$ states before being normalised (and therefore averaged). This operation is implemented by $A_{\text{POOL}}=\frac{1}{2}(I+\text{SUBTRACT}_{1})$, where $\text{SUBTRACT}_{1}\lvert x\rangle=\lvert x\ominus 1\rangle$ and $\text{ADD}_{1}\lvert x\rangle=\lvert x\oplus 1\rangle$. We can see that by taking
@@
by using the LCU framework from Section 1.4. The terms $\Omega$ and $\Omega^{\prime}$ are the normalisation constants for the initial and final states, respectively.
@@
where ()(...)( … ) collects terms with $\langle i\rvert\langle j\rvert$ such that $i+j\geq 1$ and will therefore not be used in the algorithm and can be ignored. This can be clearly implemented by a unitary operation. It can also be shown that $T_{\Delta x,\Delta y}$ can be realised with a unitary operator. Binary addition is possible to implement on a quantum circuit Vedral_1996 as shown in Figure 6 that shows $\text{ADD}_{1}\lvert x\rangle=\lvert x\oplus 1\rangle$. In general, we shall define a $k$ adding operator as $\text{ADD}_{k}\lvert x\rangle=\lvert x\oplus k\rangle$, which can be formed by applying $ADD_{1}$ in series $k$ times. The inverse of this operator, found by taking the Hermitian conjugate, will subsequently be the subtraction operator $\text{SUBTRACT}_{k}\lvert x\rangle=\lvert x\ominus k\rangle$.
Figure 6: Example circuit architecture to perform binary addition increasing all binary values by one. This takes $\text{ADD}_{1}\lvert x\rangle=\lvert x\oplus 1\rangle$ Vedral_1996 .
@@
In order to implement an average pooling layer with a pooling window size of $D\times D$ one needs to implement a linear combination $D^{2}$ unitaries via the LCU method, whereby the unitaries correspond to selecting the pixels within the tile. Due to the fact that we use two registers for $x$ and $y$, this means that we need $2D$ unitaries overall, which are controlled $\hat{T}_{\Delta x,0}$ and $\hat{T}_{0,\Delta y}$ operators acting on the separate $x$ and $y$ target registers, which are controlled by the ancilla registers. A visualisation of the average pooling layer for a $2\times 2$ pooling window is shown in Figure 5. In terms of the image boundaries, it would be possible to apply null padding to the image by introducing extra qubits in the $\lvert 0\rangle$ states. Without any padding, the quantum pooling window will treat the image as periodic and could include pixels from opposite sides of the image when located close to the image edge.
@@
It is worth mentioning that classical CNNs perform this averaging as a means of dimensionality reduction, as the total number of averages taken is usually less than the total number of pixels in the image. In the quantum case, however, it takes no additional time to average every single pixel in the image, and therefore we presented this as the most general case. Subsequently, to truly perform the quantum analogue of pooling, certain values would be discarded to reduce the dimensionality of the problem. The exact method will depend on exactly what kind of dimensionality reduction is desired, although we discuss how this could be performed in Appendix D.
3.2 Algorithm Scaling
@@
The advantage of average pooling layers is well documented in classical neural networks Galanis2022ConvolutionalNN ; gholamalinezhad2020pooling ; Zafar2022 . The method suggested in our work would allow this feature to be implemented on quantum encoded images in quantum machine learning models. The utility of this in the context of a quantum advantage will therefore be dependent on whether other quantum operations within the model can achieve some advantage over classical models. For example, recent empirical studies for amplitude encoded image data suggest that quantum models may in some circumstances display an improved robustness against adversarial attacks compared to classical methods west2023benchmarking ; west2023drastic . Indeed, theoretical guarantees on quantum robustness have recently been reported dowling2024adversarial . This is not the focus of our work, but the quantum pooling layers we introduce may be of use in further studies regarding quantum encoded image data. However, more concretely, we can show that assuming one has access to a quantum encoded image, that the quantum average pooling layer could be implemented faster than in the classical case. Specifically, the quantum parallelism allows all the pooling for every pixel to be calculated simultaneously, removing dependence on the total number of pixels in the operation.
Considering a stride value of $1$, such that the pooling window moves one pixel at a time, to perform the averaging classically, we need to average $D^{2}$ values for $N^{2}$ pixels, giving a scaling of at least $\Omega(N^{2})$ with respect to $N$ considering the amount of averaging operations required. If one has access to an amplitude encoded image, then to perform the averaging on this quantum state for all pixels requires $D^{2}$ operators in the linear combinations of unitaries, but these come from two different registers with $D$ unitary operators each. This would therefore require ancilla qubits scaling as $k=\mathcal{O}(\log(D))$ in each ancilla register. The number of qubits required for the image encoding would be $\mathcal{O}(\log(N))$ in each target register for encoding both $x$ and $y$. A key feature of this proposed method is that the $N^{2}$ scaling term associated with applying pooling classically is removed, as the nature of quantum superposition allows us to apply average pooling to every pixel in the image simultaneously. If we assume the image is provided as an amplitude encoded state and ignore for now the consideration of the probability of success inherent in the LCU method, this leads to a quadratic speedup over the classical case.
@@
$\pi_{S}=\left\lvert\frac{1}{D^{2}}\sum_{\Delta x=1}^{D}\sum_{\Delta y=1}^{D}\hat{T}_{\Delta x,\Delta y}\lvert\psi\rangle\right\rvert^{2}.$
If we had an image in which every pixel has the exact same colour value, we can effectively set $\hat{T}_{\Delta x,\Delta y}=\mathbb{I}$ and therefore $\pi_{S}=1$. We expect real-world images to have pixels that are close in colour value to other local pixels nearby on average with the exceptions occurring at the edges of subjects within the image. We therefore expect the probability to remain relatively stable overall, although one could construct adversarial example images that result in low probabilities. To gain a practical understanding of this probability scaling we considered the MNIST fashion dataset deng2012mnist , we empirically show the scaling of the probability of success with respect to $D$ and $N$ in Figure 7 and Figure 8, respectively. These results indicate that the probability of success decreases but levels off after a certain point as $D$ increases, and that there is no discernible trend when increasing $N$. These results align with our intuition for real-world images. The general structure and content of real-world images (and their local pixel similarities) remain consistent as $N$ increases; therefore, there should be no discernible decrease in $\pi_{S}$ from increasing $N$. On the other hand, increasing $D$ means considering a larger neighbourhood of pixels. Initially, as $D$ increases, the probability of encountering edges within this neighbourhood also increases, causing $\pi_{S}$ to decrease. However, beyond a certain point, further increasing $D$ adds less new information because most of the image regions are already accounted for and edges covered, so $\pi_{S}$ levels off. The exact probability will be very image dependent; there may also exist techniques to prepare images such that the probabilities are improved, which we leave as an open question for further research.
Figure 7: Average success probability of the LCU procedure for implementing an average pooling layer as the pooling window dimension $D$ is increased, while the image size is set to $N=28$. The average is taken over 100 image samples from the MNIST database deng2012mnist , with the shaded area indicating the standard deviation of the samples.
Figure 8: Average success probability of the LCU procedure for implementing an average pooling layer as the image size $N$ is increased, while the pooling window size is kept constant at $D=3$. The average is taken over 100 image samples from the MNIST database deng2012mnist , with the shaded area indicating the standard deviation of the samples.
4 Irreducible Subspace Projections

### v2

@@ III.1 Quantum Average Pooling Implementation
A classical average pooling layer consists of a pooling window of size $D\times D$ pixels that passes over an image consisting of $N\times N$ pixels, with a certain stride $K_{\text{STRIDE}}$ Rawat and Wang (2017); Sharma et al. (2018); Al-Saffar et al. (2017). We shall focus on the case where the pooling window moves across the image one pixel at a time, corresponding to a stride value $K_{\text{STRIDE}}=1$. For each position of the pooling window, all pixels within the pooling window are averaged together, and this average is output as the value of a pixel in a new image. In classical techniques depending on the size of $K_{\text{STRIDE}}$ and $D$ the number of output pixels will be some number less than $N^{2}$ and hence the pooling corresponds to reducing the dimension of the image. In the quantum technique, there is no penalty for calculating all $N^{2}$ terms as the operations are done in parallel on the quantum state, hence we shall consider this case and leave the dimensionality reduction as a subroutine that can be implemented after the averaging.
@@
so that they can be added to the average. The pixels accessed by $T_{\Delta x,\Delta y}$ define the pooling tile. For occasions in which the pooling window covers pixels outside the image, it is common to include some padding of the image Yu et al. (2023), for example, any pixels outside the $N\times N$ region can be padded by zeros, giving a padded image of dimension $(N+2D)\times(N+2D)$.
There have been many investigations focused on finding quantum parallels to CNN models, whereby measurements are taken of certain qubits such that the dimensionality of the data is reduced. For example, in some setups the pooling layer corresponds to a variational ansatz followed by a measurement of a qubit Hur et al. (2022). In Cong et al. (2019) an effective measurement was chosen corresponding to the symmetry and details of their specific problem concerning Quantum Phase Estimation. Although these popularised methods achieve the goal of pooling through reducing the dimensionality of the quantum state, they have not focused on giving a replication of an averaging pooling layer from classical CNN models for image data. In this section, we shall therefore demonstrate how LCU methods can be used to perform average pooling on quantum encoded image data. We shall utilise amplitude encoding with distinct coordinates $x$ and $y$ for the image, although any encoding for the data could be used, as long as the correct transformations are implemented with the LCU method.
@@
where $\Omega^{2}=\sum v_{i,j}^{2}$ is a normalisation constant and $v_{i,j}\in\mathbb{R}$. Once we have an image encoded into a quantum state in this manner, we can consider the type of operation required by average pooling. In a simple case, we can consider only the $x$ register for $D=2$. We will need to apply an operation that takes $\lvert x\rangle\rightarrow\frac{1}{\Omega^{\prime}}(\lvert x\rangle+\lvert x\ominus 1\rangle)$ for some normalisation constant $\Omega^{\prime}$, to perform an averaging over pixels amplitudes. This will result in the state $\lvert x\rangle$ having its amplitude transformed to the sum of the amplitudes of the $\lvert x\rangle$ and $\lvert x\oplus 1\rangle$ states before being normalised (and therefore averaged). This operation is implemented by $A_{\text{POOL}}=\frac{1}{2}(I+\text{SUBTRACT}_{1})$, where $\text{SUBTRACT}_{1}\lvert x\rangle=\lvert x\ominus 1\rangle$ and $\text{ADD}_{1}\lvert x\rangle=\lvert x\oplus 1\rangle$. These operators are implemented by decrement and increment gates. We can see that by taking
@@
by using the LCU framework from Section I.4. The terms $\Omega$ and $\Omega^{\prime}$ are the normalisation constants for the initial and final states, respectively.
@@
where ()(...)( … ) collects terms with $\langle i\rvert\langle j\rvert$ such that $i+j\geq 1$ and will therefore not be used in the algorithm and can be ignored. This can be clearly implemented by a unitary operation. It can also be shown that $T_{\Delta x,\Delta y}$ can be realised with a unitary operator. Binary addition is possible to implement on a quantum circuit Vedral et al. (1996) as shown in Figure 11 that shows $\text{ADD}_{1}\lvert x\rangle=\lvert x\oplus 1\rangle$. In general, we shall define a $k$ adding operator as $\text{ADD}_{k}\lvert x\rangle=\lvert x\oplus k\rangle$, which may be formed by applying $ADD_{1}$ in series $k$ times or through a more optimised gate specific for that value of $k$. The inverse of this operator, found by taking the Hermitian conjugate, will subsequently be the subtraction operator $\text{SUBTRACT}_{k}\lvert x\rangle=\lvert x\ominus k\rangle$.
Figure 11: Example circuit architecture to perform binary addition increasing all binary values by one. This takes $\text{ADD}_{1}\lvert x\rangle=\lvert x\oplus 1\rangle$ Vedral et al. (1996).
@@
Through this framework the average pooling operation can be implemented as is visualised for a $2\times 2$ pooling window in Figure 10.
@@
In terms of the image boundaries, it would be possible to apply null padding to the image by introducing extra qubits in the $\lvert 0\rangle$ states. Without any padding, the quantum pooling window will treat the image as periodic and could include pixels from opposite sides of the image when located close to the image edge. It is worth mentioning that classical CNNs perform this averaging as a means of dimensionality reduction, as the total number of averages taken is usually less than the total number of pixels in the image. In the quantum case, however, it takes no additional time to average every single pixel in the image, and therefore we presented this as the most general case. Subsequently, to truly perform the quantum analogue of pooling, certain values would be discarded to reduce the dimensionality of the problem. The exact method will depend on exactly what kind of dimensionality reduction is desired, although we discuss how this could be performed in Appendix D.
III.2 Algorithm Scaling
@@
The advantage of average pooling layers is well documented in classical neural networks Galanis et al. (2022); Gholamalinezhad and Khosravi (2020); Zafar et al. (2022). The method suggested in our work would allow this feature to be implemented on quantum encoded images in quantum machine learning models. The utility of this in the context of a quantum advantage will therefore be dependent on whether other quantum operations within the model can achieve some advantage over classical models. For example, recent empirical studies for amplitude encoded image data suggest that quantum models may in some circumstances display an improved robustness against adversarial attacks compared to classical methods West et al. (2023a, b). Indeed, theoretical guarantees on quantum robustness have recently been reported Dowling et al. (2024). While this is not the focus of our work, the quantum average pooling layers we introduce may be of use in further studies regarding quantum encoded image data. We will therefore focus instead on examining the circuit complexity and probabilistic scaling of the method.
III.2.1 Circuit Scaling
In order to implement an average pooling layer with a pooling window size of $D\times D$ one needs to implement a linear combination $D^{2}$ unitaries via the LCU method, whereby these unitaries correspond to selecting all pixels within the average pooling window. This means that in a given $x/y$ register one needs to create a linear combination of $D$ unitaries. A general technique for doing this using only $\log(D)$ ancilla for any LCU circuit would be to use $D$ multi-controlled operators, whereby each ancilla basis state activates exactly one of the multi-controlled operators and no others i.e. $S_{\text{SELECT}}\lvert i\rangle\lvert\psi\rangle=\lvert i\rangle\hat{T}_{i}\lvert\psi\rangle$. Previous related work on implementing general filter masks for quantum convolutional layers used this approach where they state that the number of multi-controlled unitaries that must be implemented is equal to the size of the filter mask $D^{2}$ Wei et al. (2022); Chen et al. (2022). However, we show that by focusing on average pooling and considering that subtraction operators form a closed set, that a more efficient circuit implementation is possible which uses $\mathcal{O}(\log(D))$ single-qubit controlled operators.
Without loss of generality consider only the $x$ register. We need to ensure that this has an equal superposition of $\hat{T}_{\Delta x}$ terms such that $\Delta x$ runs from 00 to $D-1$. This can be achieved using $\mathcal{O}(\log(D))$ qubits and single qubit controlled applications of $\hat{T}_{k}$ operators. In this case $\hat{T}_{k}$ is implemented by the $k$-th ancilla qubit and corresponds to the $\hat{T}_{k}=\text{SUBTRACT}_{2^{k-1}}$ operation. This operation subtracts $2^{k-1}$ from the binary register and can be formed by the $\text{SUBTRACT}_{1}$ operation applied to the $n-k+1$ most significant qubits and the identity operator applied to the $k-1$ least significant qubits. If $k\in[1,L]$, where $L$ is the total number of ancilla qubits and controlled operations then we have a set of operators $T=[\hat{T}_{1},\hat{T}_{2},\hat{T}_{4},...,\hat{T}_{2^{L-1}}]$ where the $k$-th element is applied if the $k$-th ancilla is in the 1 state. As we have an equal superposition of all $2^{L}$ basis states in the ancillas, this means that we will find an equal superposition of $2^{L}$ terms which consist of all possible combinations of the operators in $T$. The ancilla basis states can be thought of as selecting subsets of $T$, with the resultant operation found by multiplying all operators together in that subset such that the indices decrease from left to right. As we have chosen the scaling of these elements to be $\hat{T}_{k}=\text{SUBTRACT}_{2^{k-1}}$, we can see that considering all possible combinations of these terms will result in an equal sum of $\hat{T}_{\Delta x}$ operators where $\Delta x$ runs from 00 to $12^{L}-$ 1.
We can then set $D=2^{L}$ to see that the number of ancillas and controlled operators must scale as $L=\mathcal{O}(\log(D))$. Therefore, the procedure requires $\mathcal{O}(\log(D))$ unitaries that correspond to subtraction operators, that act on the separate $x$ and $y$ target registers and are controlled by $\mathcal{O}(\log(D))$ ancilla registers. An example circuit which involves 4 ancilla qubits and 4 controlled unitaries is shown in Figure 12 in which 16 total combinations of operators are implemented. Note that if $D$ is less than $2^{L}$ then the final layer could be adjusted to be $\hat{T}_{D-2^{L-1}}$ but we may require multi-control qubits in order to prevent degeneracy and maintain and equally weighted combination of operators, see Appendix F.
Figure 12: Example efficient LCU circuit which implements a linear combination of 16 operators in the range $[\mathbb{I},\hat{T}_{1},\hat{T}_{2},...,\hat{T}_{15}]$. Each operator is controlled by a single qubit and corresponds to a subtraction operator, where the amount subtracted is doubled for each successive qubit. This approach is more efficient than the most general possible implementation and is possible due to the fact that subtraction operators applied in sequence form a closed set.
To perform the averaging classically for a stride size $K$ it would require finding the average of $D^{2}$ values for $N^{2}$ pixels, giving a scaling of $\mathcal{O}(D^{2}N^{2})$ in general. Therefore classical averaging requires $\mathcal{O}(D^{2}N^{2})$ arithmetic operations, while the quantum average pooling we suggest requires $\mathcal{O}(\log(D))$ controlled subtraction operations. The exact scaling of the controlled subtraction operations depends on if ancilla qubits can be utilised, however we show this uses at worst $\mathcal{O}((\log(N))^{2})$ basic operations (see Appendix E) giving a total of $\mathcal{O}(\log(D)(\log(N))^{2})$ basic operations in the overall LCU method while using $\mathcal{O}(\log(N))$ target qubits for the image encodings and $\mathcal{O}(\log(D))$ ancilla qubits.
Previous work built an LCU based quantum convolutional layer using a $3\times 3$ filter of the form
$\begin{pmatrix}\omega_{00}\quad\omega_{01}\quad\omega_{02}\\ \omega_{10}\quad\omega_{11}\quad\omega_{12}\\ \omega_{20}\quad\omega_{21}\quad\omega_{22}\end{pmatrix}$ (79)
such that the filter passes over the image and calculates $\sum\omega_{\Delta x\Delta y}v_{i+\Delta x,j+\Delta y}$ for the pixel value in the window Wei et al. (2022). They show that average pooling can be obtained as a special case where $\omega_{ij}=1;\forall i,j$. We note that this work did not explicitly show a proof for a general $D\times D$ window, instead considering only $D=3$. They used multi-controlled operators in which each ancilla qubit state only implements exactly one unitary, meaning that their technique requires $D^{2}$ multi-controlled operators. They report their unitary operations as consisting of $\mathcal{O}((\log(N))^{6})$ basic operators, which would correspond to an overall scaling of $\mathcal{O}(D^{2}(\log(N))^{6})$. Hence, the circuit implementation that we present, which scales as $\mathcal{O}(\log(D)(\log(N))^{2})$, corresponds to a polynomial advantage in $N$ and an exponential advantage in $D$ over this previous technique when considering average pooling specifically.
Furthermore, although we started with the more specific case of average pooling, our entire algorithm can be generalised to recreate the quantum convolutional filters as presented in Wei et al. (2022). This is achieved by introducing amplitudes into the ancilla qubits instead of using an equal superposition. We see this by considering a new preparation operator in which all ancillas (for both $x$ and $y$ registers) are entangled such that
$=\frac{1}{D}\Big{(}\sum_{\Delta x=0}^{D-1}\sum_{\Delta y=0}^{D-1}\sqrt{\omega_{\Delta x,\Delta y}}\lvert\Delta x\rangle\lvert\Delta y\rangle\ \Big{)}\lvert\psi\rangle,$ (80)
where for simplicity we say $\sqrt{\omega_{\Delta x,\Delta y}}\in\mathbb{R}$ is a real amplitude of the ancilla basis state $\lvert\Delta x\rangle\lvert\Delta y\rangle$. This will mean that the terms in the linear combination of $\hat{T}_{\Delta x,\Delta y}$ operators will have an associated weighting equal to $\omega_{\Delta x,\Delta y}$. Therefore the transformation on the pixel value amplitudes in the image will be
$v_{i,j}^{\prime}=\frac{1}{D^{2}}\sum_{\Delta x=0}^{D-1}\sum_{\Delta y=0}^{D-1}\omega_{\Delta x,\Delta y}v_{i+\Delta x,j+\Delta y}.$ (81)
We can therefore identify that this performs a general quantum convolutional filter as originally described in Wei et al. (2022), while utilising exponentially fewer unitary operators in terms of $D$. A comparison of this reduction in circuit complexity is shown in Figure 13 when applying a $4\times 4$ convolutional filter defined as
$\begin{pmatrix}\omega_{00}\quad\omega_{01}\quad\omega_{02}\quad\omega_{03}\\ \omega_{10}\quad\omega_{11}\quad\omega_{12}\quad\omega_{13}\\ \omega_{20}\quad\omega_{21}\quad\omega_{22}\quad\omega_{23}\\ \omega_{30}\quad\omega_{31}\quad\omega_{32}\quad\omega_{33}\end{pmatrix}$ (82)
This general convolution case reduces to the previously discussed averaging case when the ancilla qubits are in an equal superposition. We highlight that in this section we do not discuss the actual pooling step itself, and hence any improvement is confined to the averaging / convolutional subroutine of the circuit. Further details on the practical implementation of the pooling step and quantum convolutional neural network framework can be found in Wei et al. (2022).
Figure 13: a) The circuit used to implement quantum convolutional filters in previous work Wei et al. (2022). For an equal comparison we show the generalisation of their method to allow the implementation of a $4\times 4$ size convolutional window, using $16$ multi-controlled unitaries. b) White circles correspond to controls that activate if the control qubit is in the $\lvert 0\rangle$ state. The $t$ labelled controlled-gates apply the unitary operator $Q_{t}$ which when combined using the LCU technique reconstruct the desired convolutional filter, as outlined in Wei et al. (2022). A complete definition of the $Q_{t}$ unitaries within their framework is extensive and omitted here. For both circuits $S$ is the unitary used to initialise the ancilla qubit amplitudes corresponding to the desired filter, which is then achieved when the ancillas are all measured in the $\lvert 0000\rangle$ state at the end of the circuit. c) The circuit implementation of the same $4\times 4$ convolutional filter using our framework as described in the main text. Only a single qubit controls each $\hat{T}_{\ominus t}$ which acts to subtract $t$ in binary on the $x$ or $y$ register. This arithmetic interpretation simplifies the circuit considerably as we can see each ancilla basis state applies one of $16$ different transformations $\hat{T}_{\Delta x,\Delta y}$ on the pixel registers, which provides all the transformations required to apply a $4\times 4$ size convolutional window, using only 4 controlled unitary gates in total, rather than having to implement $16$ terms individually as multi-controlled unitary gates. For example $\lvert 1111\rangle$ will be responsible for $\Delta x=3,\Delta y=3$, hence the ancilla amplitude $\alpha_{1111}$ should control the relative weight of the top right pixel of the convolution filter ($\omega_{03}$ in Equation 79), while $\alpha_{0000}$ corresponds to the identity and would control the bottom left pixel weighting ($\omega_{30}$ in Equation 79).
We also note that our implementation is easily generalised to data types with dimension greater than two by adding additional registers for extra dimensions. For $Q$-dimensional data we require $Q\log(N)$ qubits to encode the image into $Q$ different registers. In this case, the window has a total $Q$-dimensional volume of $D^{Q}$, and hence using the implementation from previous work Wei et al. (2022) would require $D^{Q}$ unitaries. In contrast, our technique would require $Q\log(D)$ operators in total with $Q\log(D)$ ancilla total qubits when considering all registers and would therefore use exponentially fewer unitaries in both data dimension $Q$ and window size $D$.
III.2.2 Probabilistic Scaling
@@
$\pi_{S}=\left\lvert\frac{1}{D^{2}}\sum_{\Delta x=0}^{D-1}\sum_{\Delta y=0}^{D-1}\hat{T}_{\Delta x,\Delta y}\lvert\psi\rangle\right\rvert^{2}.$
If we had an image in which every pixel has the exact same colour value, we can effectively set $\hat{T}_{\Delta x,\Delta y}=\mathbb{I}$ and therefore $\pi_{S}=1$. We expect real-world images to have pixels that are close in colour value to other local pixels nearby on average with the exceptions occurring at the edges of subjects within the image. We therefore expect the probability to remain relatively stable overall, although one could construct adversarial example images that result in low probabilities. To gain a practical understanding of this probability scaling we considered the MNIST fashion dataset Deng (2012), we empirically show the scaling of the probability of success with respect to $D$ and $N$ in Figure 14 and Figure 15, respectively. These results indicate that the probability of success decreases but levels off after a certain point as $D$ increases, and that there is no discernible trend when increasing $N$. These results align with our intuition for real-world images. The general structure and content of real-world images (and their local pixel similarities) remain consistent as $N$ increases; therefore, there should be no discernible decrease in $\pi_{S}$ from increasing $N$. On the other hand, increasing $D$ means considering a larger neighbourhood of pixels. Initially, as $D$ increases, the probability of encountering edges within this neighbourhood also increases, causing $\pi_{S}$ to decrease. However, beyond a certain point, further increasing $D$ adds less new information because most of the image regions are already accounted for and edges covered, so $\pi_{S}$ levels off. The exact probability will be very image dependent; there may also exist techniques to prepare images such that the probabilities are improved, which we leave as an open question for further research. In cases where the probability of success remains high, it corresponds to pixels being similar to each other locally, this in turn means that the average pooling action on those qubits locally will be well approximated by the identity operation. It is important to note therefore that these conditions could be more prone to classical simulation techniques and further work should be carried out on this possibility before any concrete advantages are claimed.
Figure 14: Average success probability of the LCU procedure for implementing an average pooling layer as the pooling window dimension $D$ is increased, while the image size is set to $N=28$. The average is taken over 100 image samples from the MNIST database Deng (2012), with the shaded area indicating the standard deviation of the samples.
Figure 15: Average success probability of the LCU procedure for implementing an average pooling layer as the image size $N$ is increased, while the pooling window size is kept constant at $D=3$. The average is taken over 100 image samples from the MNIST database Deng (2012), with the shaded area indicating the standard deviation of the samples.
IV Irreducible Subspace Projections